# Subhalos are Distributed Anisotropically About Their Hosts

### v1

@@ 6.1 Conclusion
  • Both cluster-mass and Milky Way-mass host halo’s major axes are generally well aligned with those of their subhalos. In this broad sense, subhalos are preferentially oriented in the same way as the overall mass distribution of their host halos. For Milky Way-mass halos, the median angle between the host halo major axis the subhalo component major axis is $\theta$ = $27\pm 7.7^{\circ}$ or $|\cos(\theta)|$ = 0.89 $\pm$ 0.04. For Cluster-mass halos, this median angle is $\theta$ = $15\pm 2.6^{\circ}$ or $|\cos(\theta)|$ = 0.97 $\pm$ 0.01.
  • Both cluster-mass and Milky Way-mass host halo’s angular momenta are fairly well aligned with the net angular momenta of their subhalos. For Milky Way-mass halos, the median angle between the host halo angular momentum the net subhalo angular momentum is $\theta$ = $43\pm 7.6^{\circ}$ or cos($\theta$) = 0.73 $\pm$ 0.09. For Cluster-mass halos, this median angle is $\theta$ = $56\pm 5.7^{\circ}$ or cos($\theta$) = 0.56 $\pm$ 0.08.
  • Halos that are viewed along their major axis have a greater projected subhalo mass near their centers compared to those viewed along a random axis direction. To within our ability to measure, this preference does not have a strong dependence upon the mass of the subhalo.
@@ 6.2 Discussion
There are possibly other important consequences of the alignments highlighted in this work. Our finding that subhalo and host halo angular momenta are aligned has consequences for galaxy formation studies. For example, it is common practice for semi-analytic galaxy formation models to adopt host halo angular momentum as a proxy for galaxy angular momentum either probabilistically (Somerville et al., 2008; Guo et al., 2011) or deterministically (Benson, 2012). Interestingly, our results suggest that some of the halo angular momentum is carried by subhalos that orbit in the same plane as any central disk galaxy. Sufficiently large subhalos will likely host their own galaxies. This provides an interesting contrast to a now well-known observation in the Local Group. The dwarf satellite galaxies of the Milky Way seem to orbit the Milky Way in a plane that is nearly perpendicular to the galaxy disk so that the angular momentum carried by the dwarf satellites is nearly perpendicular to that angular momentum of the Milky Way galaxy (Kroupa et al., 2005; Pawlowski, 2018). Satellites of Andromeda are found to be distributed anisotropically (Doliva-Dolinsky et al., 2023) and orbit the disk of Andromeda in a separate plane at a tilt of $\sim$50 (Pawlowski, 2018) with respect to the angular momentum of Andromeda. Such planar orbits of satellites have also been found in simulations using Milky Way and Andromeda galaxy analogs in Santos-Santos et al. (2020) and Samuel et al. (2021) with the latter showing that they can exist in a plane perpendicular to the galactic disk. However, recent work indicates that the perpendicular alignment seen in the Milky Way could be temporary and due to a chance alignment with its two most distant satellites (Leo I and II) (Sawala et al., 2023; Sales & Navarro, 2023). Additionally, prior simulation work by Fielder et al. (2018) found host halo angular momentum to be the weakest indicator of subhalo abundance, further reinforcing the idea of alignment between host halo and subhalo angular momenta. All things considered, we believe that caution must be taken by galaxy formation studies when defining galaxy angular momentum.

### v2

@@ 6.1 Conclusion
  • Both cluster-mass and Milky Way-mass host halo’s major axes are generally well aligned with those of their subhalos. In this broad sense, subhalos are preferentially oriented in the same way as the overall mass distribution of their host halos. For Milky Way-mass halos, the median angle between the host halo major axis the subhalo component major axis is $\theta$ = $30\pm 4^{\circ}$ or $|\cos(\theta)|$ = 0.86 $\pm$ 0.04. For Cluster-mass halos, this median angle is $\theta$ = $25\pm 3.5^{\circ}$ or $|\cos(\theta)|$ = 0.91 $\pm$ 0.03.
  • Both cluster-mass and Milky Way-mass host halo’s angular momenta are fairly well aligned with the net angular momenta of their subhalos. For Milky Way-mass halos, the median angle between the host halo angular momentum the net subhalo angular momentum is $\theta$ = $50\pm 9^{\circ}$ or cos($\theta$) = 0.63 $\pm$ 0.12. For Cluster-mass halos, this median angle is $\theta$ = $56\pm 5.5^{\circ}$ or cos($\theta$) = 0.56 $\pm$ 0.08.
  • Halos that are viewed along their major axis have a greater projected subhalo mass near their centers compared to those viewed along a random axis direction. As seen in the top panels of Fig. 6, within projected radii of a few percent of the virial radius of the host halo, the fraction of mass in subhalos is $\sim 175\%$ larger for Milky Way mass halos and as much as $\sim 195\%$ larger for cluster halos when projected along the major axis as compared to the average from a random projection. To within our ability to measure, this preference does not have a strong dependence upon the mass of the subhalo.
@@ 6.2 Discussion
There are possibly other important consequences of the alignments highlighted in this work. Our finding that subhalo and host halo angular momenta are aligned has consequences for galaxy formation studies. For example, it is common practice for semi-analytic galaxy formation models to adopt host halo angular momentum as a proxy for galaxy angular momentum either probabilistically (Somerville et al., 2008; Guo et al., 2011) or deterministically (Benson, 2012). Interestingly, our results suggest that a large fraction of the net angular momentum of the host halo is carried by the subhalos reside on orbits that are close to the plane of any galactic disk. Sufficiently large subhalos will likely host their own galaxies. This provides an interesting contrast to a now well-known observation in the Local Group. The dwarf satellite galaxies of the Milky Way seem to orbit the Milky Way in a plane that is nearly perpendicular to the galaxy disk so that the angular momentum carried by the dwarf satellites is nearly perpendicular to that angular momentum of the Milky Way galaxy (Kroupa et al., 2005; Pawlowski, 2018). Satellites of Andromeda are found to be distributed anisotropically (Doliva-Dolinsky et al., 2023) and orbit the disk of Andromeda in a separate plane at a tilt of $\sim$50 (Pawlowski, 2018) with respect to the angular momentum of Andromeda. Such planar orbits of satellites have also been found in simulations using Milky Way and Andromeda galaxy analogs in Santos-Santos et al. (2020) and Samuel et al. (2021) with the latter showing that they can exist in a plane perpendicular to the galactic disk. However, recent work indicates that the perpendicular alignment seen in the Milky Way could be temporary and due to a chance alignment with its two most distant satellites (Leo I and II) (Sales & Navarro, 2023). Additionally, prior simulation work by Fielder et al. (2018) found host halo angular momentum to be the weakest indicator of subhalo abundance, further reinforcing the idea of alignment between host halo and subhalo angular momenta. All things considered, we believe that caution must be taken by galaxy formation studies when defining galaxy angular momentum.